# GraphArena: Benchmarking Large Language Models on Graph Computational Problems

### v1

@@ Abstract
The “arms race” of Large Language Models (LLMs) demands novel, challenging, and diverse benchmarks to faithfully examine their progresses. We introduce GraphArena, a benchmarking tool designed to evaluate LLMs on graph computational problems using million-scale real-world graphs from diverse scenarios such as knowledge graphs, social networks, and molecular structures. GraphArena offers a suite of 10 computational tasks, encompassing four polynomial-time (e.g., Shortest Distance) and six NP-complete challenges (e.g., Travelling Salesman Problem). It features a rigorous evaluation framework that classifies LLM outputs as correct, suboptimal (feasible but not optimal), or hallucinatory (properly formatted but infeasible). Evaluation of 10 leading LLMs, including GPT-4o and LLaMA3-70B-Instruct, reveals that even top-performing models struggle with larger, more complex graph problems and exhibit hallucination issues. Despite the application of strategies such as chain-of-thought prompting, these issues remain unresolved. GraphArena contributes a valuable supplement to the existing LLM benchmarks and is open-sourced at https://github.com/squareRoot3/GraphArena.
@@ 1 Introduction
As Large Language Models (LLMs) continue to evolve, accurately assessing their expanding capabilities becomes crucial but increasingly complex. Standardized tests, which are originally designed for humans, have become common LLM benchmarks in domains such as mathematics [30, 22, 50, 16], medicine [48, 15, 40], and multi-task collections [29, 65, 74, 89]. Despite their utility, concerns about data leakage in pretraining corpora [4, 82] cast doubt on whether LLMs are genuinely reasoning or merely memorizing. Crowd-sourced evaluations [88, 20], while reflective of real-world performance, are expensive and time-consuming [13]. Alternatively, evaluations using synthetic data like theorem proving [73, 80, 32], are rigorous but often lack real-world relevance.
To balance factors like cost, novelty, complexity, diversity, and practical applicability in LLM evaluation, the community is expanding its attention to modalities that go beyond text to include vision [50, 27], audio [84], tabular [66], and spatiotemporal [49] data. Among these, graphs stand out as a particularly promising avenue. Ubiquitous in many modern applications, graphs provide a robust framework for testing LLMs’ capabilities on interpret relational information, process non-sequential data, and generalize across diverse structures[44, 36, 59, 75, 18, 19]. Specifically, graph computational problems—which typically require systematic traversal or search algorithms to solve—serve as an excellent testbed for LLMs [25, 75]. Consider the task of finding the shortest path: a model must understand the entire graph structure, identify relevant nodes, logically deduce multiple steps to trace paths between nodes, and perform calculations to arrive at a correct conclusion.
Over the past year, several studies have assessed LLM performance on graph computational problems, including notable efforts like NLGraph [75], GraphQA [25], VisionGraph [45], GITA [79], and GraphInstruct [17]. While these studies have provided valuable insights, we identify three significant shortcomings in the existing benchmarks. First, they rely on synthetic graphs generated from models such as Erdős-Rényi [24] and Barabási-Albert [5], which may not accurately reflect real-world diversity. Second, these benchmarks focus predominantly on elementary graph tasks involving small-scale graphs, limiting tasks to simple queries and counts of basic elements like nodes, edges, triangles, and rings. More complex and challenging graph problems are largely overlooked. Third, they only require LLMs to provide straightforward answers—such as yes, no, or a numerical value—rather than detailed paths or reasoning processes. This format may be utilized by models through guesswork rather than through real comprehension of the underlying logic.
@@
In this paper, we introduce GraphArena, a benchmarking tool designed to assess the reasoning capabilities of LLMs on addressing graph computational problems. Recognizing the limitations of existing benchmarks, GraphArena incorporates three core principles:
@@
  • Realistic Graph Collection. GraphArena utilizes a rich assortment of million-scale real-world graphs, covering knowledge graphs, social networks, molecular structures, etc. We aim to maintain the integrity of the original graph topology and attributes throughout the sampling process.
  • Curated Task Selection. Our benchmark offers a diverse and challenging set of tasks that examine a broad spectrum of LLM reasoning skills. Illustrated in Figure 1, GraphArena includes 4 polynomial-time and 6 NP-complete challenges, which test both intuitive (System 1) and deliberative (System 2) cognitive processes [37]. These tasks require skills from simple computations to complex multi-step analytics, long-range location identification, and heuristic approximations.
  • Rigorous Evaluation Framework. GraphArena introduces a task-specific, fine-grained evaluation protocol to increase the rigor of assessments and reduce reliance on superficial pattern recognition. LLMs are required to identify the specific component or path in the graph that leads to the solution. Responses are then classified as correct, suboptimal (feasible but not optimal), hallucinatory (properly formatted but infeasible), or missing, enabling a nuanced comparison among LLMs.
Our comprehensive assessment of ten popular LLMs across 10,000 problems in GraphArena reveals substantial insights. Consistent with common perception, GPT-4o [2] and Llama3-70b-Instruct [3] stand out as the best closed-source and open-source models, respectively. However, even these leading models face challenges with NP-complete tasks and exhibit a propensity for hallucination, in which they frequently generate structurally correct but contextually infeasible responses. The hallucination issue is more severe on larger graphs and models with fewer parameters. While chain-of-thought prompting has shown some improvements, it remains insufficient to fully resolve the hallucination issue. Additionally, graph-aware LLMs fine-tuned on relevant corpora also fail to resolve this problem. These findings indicate that graph computational problems is a valuable testbed for evaluating and advancing LLMs towards the artificial general intelligence with strong reasoning capacities.
@@ 2 Benchmark Construction
Figure 1 provides an overview of our proposed GraphArena benchmark. In this section, we first discuss the collection and sampling methodologies for real-world graphs in Section 2.1, then offers detailed descriptions of each task within GraphArena and outlines the problem generation process in Section 2.2. Finally, Section 2.3 details the evaluation framework used to score the responses.
@@ 2.1 Dataset Collection
GraphArena distinguishes itself from previous benchmarks by utilizing real-world graphs instead of synthetic ones, offering a richer diversity of data. We collect graphs from the following five sources:
DBLP [39] is an academic database containing over 7 million publications under the CC0 1.0 Public Domain Dedication license. We use an undirected research collaboration graph for the period 2003-2018 from DBLP [72]. This graph consists of 1,354,852 nodes and 5,129,047 edges, with each node representing an author and each edge reflecting a collaboration that has occurred at least five times.
Social Network [63] is a repository comprising over 70 different social networks. We combine graphs from four large networks—Digg, Flixster, Lastfm, and Pokec—to form a dataset of 6,118,793 nodes and 40,647,227 edges. Each node symbolizes a user, and each edge denotes a friendship connection. All user names are pseudonymized for privacy. It is under the CC BY-SA 3.0 License.
DBpedia [11] extracts structured content from Wikipedia to form a comprehensive knowledge graph. We use DBpedia1M [81], a subset of the English version, which comprises 1,160,306 nodes and 7,214,631 edges. Each node represents an entity, and each edge denotes a relationship, with all original textual content retained as attributes. It is under the CC BY-SA 3.0 License.
Openflights [55] provides data for over 10,000 airports and their connecting flight routes. We construct an undirected graph where nodes are airports and edges are flight routes, weighted by the geographical distance between airports. We preserve the largest connected component in this graph, which consists of 3,390 nodes and 19,166 edges. It is under the DbCL v1.0 License.
PubChemQC [53] offers a quantum chemistry database detailing ground-state electronic structures. From the PCQM4Mv2 Dataset—part of the OGB Large-Scale Challenge [31] collected from PubChemQC—we incorporate 3,746,620 individual graphs. Each graph represents a molecule, where nodes are atoms and edges are chemical bonds, with atom types preserved as node attributes. It is under the CC BY 4.0 License.
We employ a random walk with restart strategy initiating from randomly selected nodes to effectively sample subgraphs. This technique enables us to identify the top-k most frequently visited nodes, forming local dense subgraphs that closely maintain the topological features of the original graphs. For the PubChemQC dataset, we sample molecular graphs directly by specified sizes. All graphs are treated as unweighted and undirected, except for OpenFlights which is weighted.
2.2 Task Design
Wikipedia lists 73 graph computational problems in graph theory, not counting variants. Selecting an appropriate subset for benchmarking is a notable challenge. Previous benchmarks often focused on simpler tasks like querying the existence and counts of basic graph elements such as nodes, edges, triangles, and rings [75, 25]. In contrast, we aim at incorporating more complex problems that demand advanced reasoning abilities.
To establish a nuanced evaluation of responses, we decompose each task into two requirements: satisfying the first requirement indicates a feasible solution, whereas fulfilling both requirements denotes a correct solution. We have chosen four polynomial-time tasks that require increasingly thorough understandings of graph structures:
@@ 2.2 Task Design
Each task is briefly described by a keyword (Neighbor, Distance, Component, Diameter) to facilitate easy reference. For the NP-complete challenges, we have selected six representative tasks:

### v2

# GraphArena: Evaluating and Exploring Large Language Models on Graph Computation

@@ Abstract
The “arms race” of Large Language Models (LLMs) demands new benchmarks to examine their progresses. In this paper, we introduce GraphArena, a benchmarking tool designed to evaluate LLMs on real-world graph computational problems. It offers a suite of four polynomial-time tasks (e.g., Shortest Distance) and six NP-complete challenges (e.g., Traveling Salesman Problem). GraphArena features a rigorous evaluation framework that classifies LLM outputs as correct, suboptimal (feasible but not optimal), hallucinatory (properly formatted but infeasible), or missing. Evaluation of over 10 LLMs reveals that even top-performing LLMs struggle with larger, more complex graph problems and exhibit hallucination issues. We further explore four potential solutions to address this issue and improve LLMs on graph computation, including chain-of-thought prompting, instruction tuning, code writing, and scaling test-time compute, each demonstrating unique strengths and limitations. GraphArena complements the existing LLM benchmarks and is open-sourced at https://github.com/squareRoot3/GraphArena.
@@ 1 Introduction
Evaluating LLMs, especially their advanced reasoning capabilities, remains a significant long-term challenge. While standardized tests in fields like mathematics Hendrycks et al. (2021); Cobbe et al. (2021), medicine Chen et al. (2023); Li et al. (2023a), and multi-task collections Hendrycks et al. (2020); Srivastava et al. (2022) have become popular LLM benchmarks, concerns about data leakage in pretraining corpora Aiyappa et al. (2023); Xu et al. (2024) raise questions about whether LLMs are genuinely reasoning or merely memorizing information. Crowd-sourced evaluations can address this issue but are typically expensive and time-consuming Zheng et al. (2023b); Chiang et al. (2024); Chang et al. (2024). Alternatively, evaluations using synthetic data, such as theorem proving Trinh et al. (2024); Xin et al. (2024), offer more rigorous assessments but often lack real-world relevance.
To balance factors such as cost, novelty, complexity, diversity, and practical applicability in LLM evaluation, the research community has expanded its focus beyond text to include various modalities, such as vision Lu et al. (2024), tabular Sui et al. (2024), and spatiotemporal data Lopez-Lira & Tang (2023). Among these modalities, graphs emerge as a particularly promising avenue for evaluating LLMs’ capabilities in several key areas: interpreting relational information or structured knowledge, processing non-sequential or non-Euclidean data, and generalizing across diverse structures or tasks. The potential of graph-based evaluations for LLMs has been recognized and explored in recent literature Perozzi et al. (2024); Wang et al. (2023); Chen et al. (2024b).
In this paper, we benchmark LLMs on graph computational problems, which often require systematic traversal or search algorithms to solve. These problems serve as an ideal testbed for evaluating the reasoning capabilities of LLMs Li et al. (2024); Wei et al. (2024); Chen et al. (2024a). While several graph problem-solving benchmarks have been established, such as NLGraph Wang et al. (2023) and GraphQA Fatemi et al. (2023), as well as algorithmic reasoning datasets like CLRS-Text Markeeva et al. (2024) and MAGMA Taylor et al. (2024), we identify several limitations in the existing landscape. First, these datasets are predominantly built on synthetic graphs, such as those generated by the Erdős-Rényi ERDdS & R&wi (1959) models, which may not adequately reflect real-world diversity. Second, the tasks within these benchmarks are generally confined to basic structural understanding (e.g., connectivity checks) and algorithm execution (e.g., breadth-first search). They largely overlook the assessment of higher-order reasoning skills, such as problem abstraction and simplification, strategy comparision and selection, etc. These capabilities are better demonstrated through solving more complex and open-ended tasks, such as NP-complete problems. Third, current evaluation methods typically rely on string matching between the final answers and the responses, allowing models to succeed through guesswork rather than demonstrating a genuine understanding of the underlying logic. Such methods lack a nuanced categorization of failure modes, limiting the depth of insights derived from the metrics.
@@
To address these issues, we introduce GraphArena in Section 2, a benchmarking tool designed to assess the reasoning capabilities of LLMs when tackling graph computational problems. Compared with previous graph benchmarks, three core improvements are introduced:
@@
  • Realistic Graph Collection. Each problem in GraphArena features a subgraph sampled from a rich assortment of real-world graphs, including knowledge graphs, social networks, and molecular structures. Our sampling process is designed to preserve the original graph topology and attributes to capture real-world diversity. Additionally, each problem is contextualized within the real-world setting of the graph, offering a more authentic and challenging evaluation environment.
  • Comprehensive Task Selection. Our benchmark presents more demanding graph tasks to assess multi-dimensional reasoning skills. As shown in Figure 1, it encompasses four polynomial-time challenges that test direct algorithmic reasoning, requiring models to accurately execute algorithm procedures step-by-step. Moreover, it includes six NP-complete task that demand meta algorithmic planning, where models must analyze problem characteristics and select suitable algorithms to excute. This selection of diverse and well-known tasks ensures a thorough evaluation of LLMs’ abilities to perform both straightforward algorithm execution and strategic decision-making.
  • Rigorous Path-based Evaluation. We propose a detailed evaluation protocol for each task to differentiate genuine problem-solving from pattern-based guessing. LLMs must generate the entire solution path (or its components) within the graph that leads to the solution. Through a three-step process involving path extraction, feasibility check, and optimality verification, responses are categorized as correct, suboptimal (feasible but not optimal), hallucinatory (properly formatted yet infeasible), or missing. This approach allows for a nuanced and accurate comparison among LLMs.
In our experiments, we extend the comparative analysis beyond LLMs by incorporating a diverse set of baseline methods, including classical graph algorithms, graph neural networks, and Graph-LLM hybrid approaches. Additionally, we explore strategies beyond conventional prompt engineering to improve LLMs’ performance in graph reasoning tasks. Our study seeks to address two key research questions:
To what degree can current LLMs solve graph computational problems? (Section 3.1) After evaluating ten popular LLMs on 10,000 GraphArena problems, we find that LLMs perform significantly better on polynomial-time tasks than on NP-complete ones. They excel in direct algorithmic reasoning but struggle with meta algorithmic planning, which demands higher-level strategic thinking. Even the most advanced LLMs, such as GPT-4o and Claude-3.5-Sonnet, struggle with graph computations—often producing hallucinatory outputs—particularly when handling larger graphs, more complex tasks, particularly with larger graphs and more complex tasks. For NP problems, these models typically default to greedy algorithms, though they sometimes utilize more effective approximation techniques. Despite their current limitations, our results highlight the potential of LLMs to serve as alternative heuristics for NP tasks, complementing existing approximation methods.
How can we further enhance LLMs on graph reasoning? (Section 3.2) While previous benchmarks for graph problems have primarily focused on prompt engineering, we observe that chain-of-thought prompting exhibits limited effectiveness on GraphArena. To overcome this limitation, we explore several advanced approaches: (1) Instruction tuning on similar graph problems and solution paths enhances performance on small-scale polynomial tasks. However, it proves less effective for large-scale NP problems, as fine-tuning alone is insufficient to develop complex reasoning abilities, such as algorithmic planning and long-range reasoning. (2) Code generation and execution by LLMs show significant promise for solving large graphs and tackling more complex tasks, though some performance degradation is observed on small graphs. (3) Increasing test-time compute and allowing LLMs multiple attempts yield modest but consistent performance improvements. These findings underscore the need for innovative strategies to advance LLMs’ graph reasoning capabilities.
@@ 2 Benchmark Construction
Figure 1 provides an overview of our proposed GraphArena benchmark. In this section, we first discuss the collection and sampling methodologies for real-world graphs in Section 2.1, then offers detailed descriptions of 10 tasks within GraphArena and the problem generation process in Section 2.2. Finally, Section 2.3 describes the evaluation framework to score and compare LLMs’ responses.
@@ 2.1 Dataset Collection
GraphArena distinguishes itself from previous benchmarks by utilizing real-world graphs, offering richer diversity compared with synthetic ones. Graphs are collected from five sources: DBLP Ley (2002), an academic collaboration network with over 1.3 million nodes (authors) and 5 million edges (collaborations); Social Network Rossi & Ahmed (2015), a composite graph combining Digg, Flixster, Lastfm, and Pokec data, totaling over 6 million nodes (users) and 40 million edges (friendships); DBpedia Bizer et al. (2009), a knowledge graph subset (DBpedia1M Xin et al. (2022)) with over 1 million nodes (entities) and 7 million edges (relationships); OpenFlights OpenFlights , an airport network consisting of 3,390 nodes (airports) and 19,166 edges (flight routes) weighted by geographical distance; and PubChemQC Nakata & Shimazaki (2017), containing 3.7 million molecular graphs from the PCQM4Mv2 Dataset Hu et al. (2021), where nodes represent atoms and edges represent chemical bonds. The detailed description of these datasets is in Appendix A.
We employ a random walk with restart strategy initiating from randomly selected nodes to sample subgraphs effectively. This enables us to identify the top-$k$ most frequently visited nodes, forming local dense subgraphs that closely maintain the topological features of the original graphs. For the PubChemQC dataset, we sample molecular graphs directly by specified sizes. All graphs are treated as unweighted and undirected, except for OpenFlights which is weighted.
2.2 Task Selection
While there is no gold standard for task selection, we focused on well-known tasks with rich real-world applications, and challenging tasks that assess multi-dimensional reasoning skills. We consider four polynomial-time tasks to test direct algorithmic reasoning, requiring models to accurately execute established algorithm procedures step-by-step:
@@ 2.2 Task Selection
Each task is briefly described by a keyword (Neighbor, Distance, Component, Diameter) to facilitate easy reference. We decompose each task into two requirements, the first basic requirement and the second optimal requirement. We further consider six NP-complete tasks to evaluate meta algorithmic planning, where models must analyze problem characteristics and select suitable algorithms to excute: